# Does Data Contamination Detection Work (Well) for LLMs? A Survey and Evaluation on Detection Assumptions

### v1

@@ Abstract
Large language models (LLMs) have demonstrated great performance across various benchmarks, showing potential as general-purpose task solvers. However, as LLMs are typically trained on vast amounts of data, a significant concern in their evaluation is data contamination, where overlap between training data and evaluation datasets inflates performance assessments. While multiple approaches have been developed to identify data contamination, these approaches rely on specific assumptions that may not hold universally across different settings. To bridge this gap, we systematically review 47 papers on data contamination detection, categorize the underlying assumptions, and assess whether they have been rigorously validated. We identify and analyze eight categories of assumptions and test three of them as case studies. Our analysis reveals that when classifying instances used for pretraining LLMs, detection approaches based on these three assumptions perform close to random guessing, suggesting that current LLMs learn data distributions rather than memorizing individual instances. Overall, this work underscores the importance of approaches clearly stating their underlying assumptions and testing their validity across various scenarios [^1] We will release the data and code associated with this work upon acceptance of this work..
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance across various benchmarks, signaling their potential to revolutionize numerous technical domains as general-purpose problem solvers Achiam et al. (2023); Meta AI (2024). However, a key concern in accurately evaluating those LLMs is the possibility of data contamination, where the LLM’s training data overlaps with the evaluation dataset Balloccu et al. (2024). Evaluating LLMs on contaminated benchmarks leads to inflated performance assessments Balloccu et al. (2024); Sainz et al. (2023a); Li and Flanigan (2024), and creates a misleading perception of their capabilities. Therefore, multiple detection approaches have been developed to identify data contamination in LLMs, and these approaches can also be deployed to identify the use of copyrighted or sensitive content in LLM training Xu et al. (2024).
@@
To bridge this gap, we (1) systematically review 47 papers on data contamination detection for LMs, (2) present the formal, mathematical definitions for different levels of data contamination, (3) categorize the underlying requirements and assumptions associated with each approach and critically assess whether these assumptions have been rigorously validated, and (4) demonstrate that some unverified assumptions can be wrong in multiple scenarios through three case studies.
@@ 2 Literature Evaluation
To systematically investigate approaches for data contamination detection, we implement a three-step literature review process with 3 sources of papers: (1) three key survey papers Xu et al. (2024); Ishihara (2023); Hu et al. (2022) and papers accepted in the 1st Workshop on Data Contamination at ACL 2024333https://conda-workshop.GitHub.io/. Three papers from this workshop were excluded from the analysis as they were not publicly available during our research.; (2) relevant papers cited by the papers from Source (1); and (3) relevant papers cited by the papers from Source (2).
From the above three sources, we collect 78 relevant papers on data contamination. We additionally filter these papers according to the inclusion criterion: the paper must propose and/or evaluate detection approaches for data contamination in text datasets and LMs. 8 studies solely discussing risks and mitigation strategies for data contamination do not meet the inclusion criteria and are excluded.
Consequently, our review includes a total of 47 papers. Among them, we systematically summarize those detection approaches and present formal mathematical representations for their underlying requirements and assumptions. We then evaluate whether these underlying assumptions are true under different scenarios, as described below.
@@ 3.1 Instance-Level Contamination
In this study, we focus on text datasets and define a language instance $x$ as a sequence of word tokens. Originally, direct instance-level contamination is defined as the presence of an instance $x$ within an LM $M$’s train set, $D_{M}$, i.e. $x\in D_{M}$ Xu et al. (2024). However, LMs often do not publish their exact train corpus, but instead refer to multiple datasets, as subsets of $D_{M}$ Zhao et al. (2023). These datasets typically undergo various pre-processing steps, such as de-duplication, filtering, masking, and removing noise. Consequently, LMs are trained on slightly different versions of the same dataset Palavalli et al. (2024). Meanwhile, there is also indirect instance-level contamination from greater variations of the dataset, such as machine paraphrasing Yang et al. (2023).
To account for such minor differences and indirect contamination, in our Definition D1 below, we introduce a Binary Indicator Function for Instance-Level Contamination, $b(x,x^{\prime})$, which outputs True (1) if two instances are considered to be the same and False (0) otherwise. Researchers can determine what instances are considered to be the same by defining $b(x,x^{\prime})$ accordingly.
@@
We define an instance $x$ to be seen by $M$, or $M$ is contaminated by $x$, iff $f(M,x)=1$. Conversely, we define an instance $x$ as clean or unseen by $M$, iff $f(M,x)=0$.
@@
Li (2023b), Riddell et al. (2024),
Deng et al. (2024), Yang et al. (2023),
@@
Table 1: Existing detection approaches for direct data contamination, their requirements and assumptions, and Critiques they received. Some papers cover multiple detection approaches with different assumptions. Most detection methods apply to both instance- and dataset-level contamination, while * denotes those limited to dataset-level contamination. Our case studies show that the underlined requirements/assumptions are not always satisfied.
@@ 3.2 Dataset-Level Contamination
When creating benchmarks for detecting data contamination, previous work typically generates the fully contaminated split. For example, Maini et al. (2024b) created contaminated and clean datasets, respectively from the train and validation splits of the LM’s pretraining corpus. Shi et al. (2023) focused on LMs which disclosed that they used Wikipedia event data for training, and created the contaminated dataset from the Wikipedia event data which were published before the LMs’ release.
@@ 4.1 Instance Similarity
Previous research focuses on developing a better or more efficient similarity function. Examples of similarity calculation can be conducted through exact match Dodge et al. (2021), fuzzy match Piktus et al. (2023); Lee et al. , automatic NLG evaluation metrics Elangovan et al. (2021); Deng et al. (2024), and another LM Yang et al. (2023). Tools have also been developed to allow efficient comparison to a large $D_{M}$, such as Data Portraits Marone and Van Durme (2023) and ROOTS Search Tool Piktus et al. (2023).
@@
Case Study: To examine how often these two requirements are met, we analyzed the top 10 LMs on the Vellum LLM leaderboard  [^4] https://www.vellum.ai/llm-leaderboard#model-comparison. Accessed on Oct 6, 2024.. We found that none of the LMs fulfilled the most basic, requirement R1, let alone R2 (see Appendix A.3 for details).
@@ 4.2 Probability Analysis
When the training dataset $D_{M}$ is unavailable, but the LM $M$’s output token probabilities are known, probability analysis has been used to detect potential instance-level contamination. We group those detection approaches by their assumptions, and unless otherwise specified, they have no requirements.
@@ 4.2.1 Absolute Probability
Seen instances will have a higher probability than unseen ones, and there exists a threshold, $\xi_{p}$, that separates seen instances from unseen ones:
@@
Previous research measures $P_{M}(x)$ through perplexity Carlini et al. (2021); Li (2023a) or approximates it through LM loss Wei et al. (2024), which can be impacted by the instance domain and simplicity. To improve upon this assumption, Shi et al. (2023) only evaluates the average probabilities of top $p\%$ least likely tokens in an instance (Min TOP $p$ Prob), assuming that unseen instances contain more low-probability outliers in Wikipedia events and books. Similarly, Song and Shmatikov (2019) assesses probabilities of the $k$ most frequent tokens.
Likewise, Srivastava et al. (2023) and Wei et al. (2024) proposed inserting special strings as watermarks into the training data, using the probability of these watermarks to detect data contamination.
@@ 4.2.2 Reference Probability by An Instance
There exists an algorithm which, given an instance $x$ and an LM $M$, can automatically generate a similar instance, unseen $x^{\prime}$.
@@
Another study, Oren et al. (2024), constructs the reference instance by randomly shuffling (exchanging) the order of sentences. They make another assumption of the exchangeability, positing that all orderings of an exchangeable benchmark should be equally likely if uncontaminated. This assumption might not be valid for coding and reasoning tasks.
4.2.3 Reference Probability by An LM
@@
Given an instance $x$, we can find another LM $M^{\prime}$, such that $x$ is unseen by $M^{\prime}$.
@@
Previous research has utilized the zlib entropy Carlini et al. (2021) and another LM Carlini et al. (2021); Mireshghallah et al. (2022) as the reference model. However, Maini et al. (2024b) and Duan et al. (2024) have demonstrated that those reference models perform close to random guessing across various domains. Cao et al. (2024) also show the zlib entropy does not work for code generation tasks.
Similarly to this assumption at the sentence level, Zanella-Béguelin et al. (2020) deploy a reference model for both individual token probability and its probability rank within the vocabulary.
@@ 4.3 Instance Generation and Instance Selection
Instance generation detects contamination by treating $x$ as a prefix-suffix pair, $x=(x_{p},x_{s})$. These approaches evaluate the LM $M$’s generated output, $M(x_{p})$, conditioned on $x_{p}$. If $M(x_{p})$ is similar or identical to $x_{s}$, $x$ will be predicted as seen. Based on this core intuition, instance generation approaches vary in their assumptions regarding input-output pairs and language generation approaches. Unless otherwise mentioned, those approaches below focus on instance generation.
@@ 4.3.1 Verbatim Memorization
Given an prefix-suffix pair $x=(x_{p},x_{s})$, if $x$ has been seen by an LM $M$, $x_{s}$ can be generated (memorized) by $M$, with the input $x_{p}$, through greedy decoding.
@@
Instance-level contamination does not always lead to verbatim memorization. Utilizing instance generation, Kandpal et al. (2022), Carlini et al. (2019), Carlini et al. (2022), and Tirumala et al. (2022) demonstrate that verbatim memorization requires repeated exposures to this instance $x$ during training, and a larger LM and longer input length $x_{p}$ can result in better memorization. Schwarzschild et al. (2024) used the minimum length of $x_{p}$ needed to generate the desired output $x_{s}$, to define the degree of memorization.
@@
For instance selection approaches, Duarte et al. (2024) and Golchin and Surdeanu (2023a) define $x_{s}$ as sentences or passages, and create similar instances to $x_{s}$ by paraphrasing $x_{s}$ using another LM. They assume that the contaminated LM will be more likely to select the verbatim option.
@@ 4.3.2 Key Information Generation
Leveraging this assumption, prior studies have masked key information within specific datasets, including input questions in NLP benchmarks Deng et al. (2024); Golchin and Surdeanu (2023b); Liu et al. (2024), column names in SQL code generation questions Ranaldi et al. (2024), character names in books Chang et al. (2023), and labels in NLI and SST tasks Magar and Schwartz (2022).
@@ 4.3.3 Generation Variation
This type of approach investigates how an LM’s outputs vary if it has seen an instance.
@@
$\text{Var}(\{M_{\cdot}(x_{p})\})\left\{\begin{array}[]{ll}\geq\xi_{v}&\text{if}f(M,x_{p})=1\\ <\xi_{v}&\text{if }f(M,x_{p})=0\end{array}\right.$ (7)
where $Var(\{M_{\cdot}(x_{p})\})$ measures the variations of outputs for $M$ under diverse, different sampling strategies when given $x_{p}$; $\xi_{v}$ is a threshold, based on the $Var(\cdot)$ and the type of input $x_{p}$.
@@ 4.3.4 Metadata-based Memorization
$\left\{\begin{array}[]{ll}\forall x\in D,S(M(x_{m}),x)<\tau_{m}&\text{if $D$ is unseen}\\ \exists x\in D,S(M(x_{m}),x)\geq\tau_{m}&\text{Otherwise}\end{array}\right.$ (8)
Here, $M(x_{m})$ might generate multiple instances when given $D$’s metadata $m$; $S(M(x_{m}),x)$ represents the highest similarity between $x$ and an instance $x^{\prime}$ as a subsequence of $M(x_{m})$; $\tau_{m}$ is the contamination threshold for $S(\cdot,\cdot)$.
@@ 4.4 Answer Memorization
Answer memorization is usually conducted at the dataset level. It introduces perturbations to the original dataset, measures the LM’s performance change, and aims to detect if the LM’s high performance is due to memorizing its answer.
@@
Given an LM $M$ and a evaluation dataset $D$, one can generate a similar dataset for the same task, $D^{\prime}$, by modifying every $x\in D$.
@@
$\text{Eval}(M,D)\left\{\begin{array}[]{ll}\not\gg\text{Eval}(M,D^{\prime})&\text{if $D$ is unseen}\\ \gg\text{Eval}(M,D^{\prime})&\text{Otherwise}\end{array}\right.$ (9)

### v3

@@ Abstract
Large language models (LLMs) have demonstrated great performance across various benchmarks, showing potential as general-purpose task solvers. However, as LLMs are typically trained on vast amounts of data, a significant concern in their evaluation is data contamination, where overlap between training data and evaluation datasets inflates performance assessments. Multiple approaches have been developed to identify data contamination. These approaches rely on specific assumptions that may not hold universally across different settings. To bridge this gap, we systematically review 50 papers on data contamination detection, categorize the underlying assumptions, and assess whether they have been rigorously validated. We identify and analyze eight categories of assumptions and test three of them as case studies. Our case studies focus on detecting direct, instance-level data contamination, which is also referred to as Membership Inference Attacks (MIA). Our analysis reveals that MIA approaches based on these three assumptions can have similar performance to random guessing, on datasets used in LLM pretraining, suggesting that current LLMs might learn data distributions rather than memorizing individual instances. Meanwhile, MIA can easily fail when there are data distribution shifts between the seen and unseen instances  [^1] Links to all relevant papers and the code for the case study are available on our project GitHub: https://github.com/velvinnn/LLM_MIA..
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance across various benchmarks, signaling their potential to revolutionize numerous technical domains as general-purpose problem solvers Achiam et al. (2023); Meta AI (2024). However, a key concern in accurately evaluating those LLMs is the possibility of data contamination, where the LLM’s training data overlaps with the evaluation dataset Balloccu et al. (2024). Evaluating LLMs on contaminated benchmarks leads to inflated performance assessments Balloccu et al. (2024); Sainz et al. (2023a); Li and Flanigan (2024), and creates a misleading perception of their capabilities. Therefore, multiple detection approaches have been developed to identify data contamination in LLMs, and these approaches can also be deployed to identify the use of copyrighted or sensitive content in LLM training Xu et al. (2024); Meeus et al. (2024b).
@@
To bridge this gap, we (1) systematically review 50 papers on data contamination detection for LMs, (2) present the formal, mathematical definitions for different levels of data contamination, (3) categorize the underlying requirements and assumptions associated with each approach and critically assess whether these assumptions have been rigorously validated, and (4) demonstrate through case studies that some unverified assumptions can be wrong in multiple scenarios.
@@ 2 Literature Evaluation
To systematically investigate approaches for data contamination detection, we implement a three-step literature review process with 3 sources of papers: (1) four key survey papers Xu et al. (2024); Ishihara (2023); Hu et al. (2022); Deng et al. (2024a) and papers from the 1st Workshop on Data Contamination at ACL 2024333https://conda-workshop.GitHub.io/.; (2) relevant papers cited by the papers from Source (1); and (3) relevant papers cited by the papers from Source (2).
From the above three sources, we collect 81 relevant papers on data contamination. We additionally filter these papers according to the inclusion criterion: the paper must propose and/or evaluate detection approaches for data contamination in text datasets and LMs. Further more, eight studies solely discussing risks and mitigation strategies for data contamination do not meet the inclusion criteria and are excluded.
Consequently, our review includes a total of 50 papers. Among them, we systematically summarize those detection approaches and present formal mathematical representations for their underlying requirements and assumptions. We then evaluate whether these underlying assumptions are true under different scenarios, as described below.
@@ 3.1 Instance-Level Contamination
In this study, we focus on text datasets and define a language instance $x$ as a sequence of word tokens. Originally, direct instance-level contamination is defined as the presence of an instance $x$ within an LM $M$’s training set, $D_{M}$, i.e. $x\in D_{M}$ Xu et al. (2024). However, LMs often do not publish their exact training corpus, but instead refer to multiple datasets, as subsets of $D_{M}$ Zhao et al. (2023). These datasets typically undergo various pre-processing steps, such as de-duplication, filtering, masking, and removing noise. Consequently, LMs are trained on slightly different versions of the same dataset Palavalli et al. (2024). Meanwhile, there is also indirect instance-level contamination from greater variations of the dataset, such as machine paraphrasing Yang et al. (2023).
To account for such minor differences and indirect contamination, in our Definition D1 below, we introduce a Binary Indicator Function for Instance-Level Contamination, $b(x,x^{\prime})$, which returns True (1) if two instances are considered to be the same and False (0) otherwise. Researchers can determine what instances are considered to be the same by defining $b(x,x^{\prime})$ accordingly.
@@
We define an instance $x$ to be seen by $M$, or $M$ is contaminated by $x$, iff $f(M,x)=1$. Conversely, we define an instance $x$ as clean or unseen by $M$ iff $f(M,x)=0$.
@@
Li et al. (2024), Riddell et al. (2024),
Deng et al. (2024b), Yang et al. (2023),
@@
Table 1: Existing detection approaches for direct data contamination, their requirements and assumptions, and critiques they received. Some papers cover multiple detection approaches with different assumptions. Most detection methods apply to both instance- and dataset-level contamination, while * denotes those limited to dataset-level contamination. In this study, we show that the underlined assumptions may not be often satisfied.
@@ 3.2 Dataset-Level Contamination
When creating benchmarks for detecting data contamination, previous work typically generates the fully contaminated split. For example, Maini et al. (2024b) created contaminated and clean datasets, respectively from the training and validation splits of the LM’s pretraining corpus. Shi et al. (2023) focused on LMs which disclosed that they used Wikipedia event data for training, and created the contaminated dataset from the Wikipedia event data which were published before the LMs’ release.
@@ 4.1 Instance Similarity
Previous research focuses on developing a better or more efficient similarity function. Examples of similarity calculation can be conducted through exact match Dodge et al. (2021), fuzzy match Piktus et al. (2023); Lee et al. , automatic NLG evaluation metrics Elangovan et al. (2021); Deng et al. (2024b), and another LM Yang et al. (2023). Tools have also been developed to allow efficient search within a large $D_{M}$, such as Data Portraits Marone and Van Durme (2023) and ROOTS Search Tool Piktus et al. (2023).
@@
Case Study: To examine how often these two requirements are met, we analyzed the top 10 LMs on the Vellum LLM leaderboard  [^4] https://www.vellum.ai/llm-leaderboard#model-comparison. Accessed on Oct 6, 2024.. We found that none of the LMs fulfilled R1, the most basic requirement, let alone R2 (see Appendix A.3 for details).
@@ 4.2 Probability Analysis
When the training dataset $D_{M}$ is unavailable, but the LM $M$’s output token probabilities are known, probability analysis has been used to detect potential instance-level contamination. We group those detection approaches by their assumptions, and unless specified otherwise, they have no requirements.
@@ 4.2.1 Absolute Probability
Seen instances will have higher probabilities than unseen ones, and there exists a threshold, $\xi_{p}$, that separates seen instances from unseen ones:
@@
Previous research measures $P_{M}(x)$ through perplexity Carlini et al. (2021); Li (2023) or approximates it through LM loss Wei et al. (2024), which can be impacted by the instance domain and simplicity. To improve upon this assumption, Shi et al. (2023) evaluates only the average token probability of top $p\%$ least likely tokens in an instance (Min $p$Token), assuming that unseen instances contain more low-probability outliers in Wikipedia events and books. Similarly, Song and Shmatikov (2019) assesses probabilities of the $k$ most frequent tokens.
Likewise, Srivastava et al. (2023), Wei et al. (2024), and Meeus et al. (2024b) proposed inserting special strings as watermarks into the training data, using the probability of these watermarks to detect data contamination.
@@ 4.2.2 Reference Probability by An Instance
There exists an algorithm which, given an instance $x$ and an LM $M$, can automatically generate a similar unseen instance, $x^{\prime}$.
@@
Another study, Oren et al. (2024), constructs the reference instance by randomly shuffling (exchanging) the order of sentences in the original instance. They make another assumption of the exchangeability, positing that all orderings of all the instances in an exchangeable benchmark should be equally likely if uncontaminated. This assumption might not be valid for coding and reasoning tasks.
4.2.3 Reference Probability by Another LM
@@
Given an instance $x$, we can find another LM $M^{\prime}$ such that $x$ is unseen by $M^{\prime}$.
@@
Previous research has utilized term frequency Meeus et al. (2024a), the zlib entropy Carlini et al. (2021), and another LM Carlini et al. (2021); Mireshghallah et al. (2022) as the reference model. However, Maini et al. (2024b) and Duan et al. (2024) have demonstrated that those reference models perform close to random guessing across various domains. Cao et al. (2024) also show the zlib entropy does not work for code generation tasks. Instead of using reference probabilities at the sentence level, Zanella-Béguelin et al. (2020) deploy a reference model for both individual token probability and its probability rank within the vocabulary.
@@ 4.3 Instance Generation and Instance Selection
Instance generation detects contamination by treating $x$ as a prefix-suffix pair, $x=(x_{p},x_{s})$. These approaches evaluate the LM $M$’s generated output, $M(x_{p})$, conditioned on $x_{p}$. If $M(x_{p})$ is similar or identical to $x_{s}$, $x$ will be predicted as seen. Based on this core intuition, instance generation approaches vary in their assumptions regarding input-output pairs and language generation approaches. Unless specified otherwise, those approaches below focus on instance generation.
@@ 4.3.1 Verbatim Memorization
Given an prefix-suffix pair $x=(x_{p},x_{s})$, if $x$ has been seen by an LM $M$, $x_{s}$ can be generated (memorized) by $M$ through greedy decoding, when given the input $x_{p}$.
@@
Duarte et al. (2024) and Golchin and Surdeanu (2023a) define $x_{s}$ as sentences or passages, and create similar instances to $x_{s}$ by paraphrasing $x_{s}$ using another LM. They use instance selection, and assume that the contaminated LM will be more likely to select the verbatim option.
However, instance-level contamination does not always lead to verbatim memorization. Utilizing instance generation, Kandpal et al. (2022), Carlini et al. (2019), Carlini et al. (2022), and Tirumala et al. (2022) demonstrate that verbatim memorization requires repeated exposures to this instance $x$ during training, and a larger LM and longer input length $x_{p}$ can result in better memorization. Schwarzschild et al. (2024) used the minimum length of $x_{p}$ needed to generate the desired output $x_{s}$ to define the degree of memorization.
@@ 4.3.2 Key Information Generation
Leveraging this assumption, prior studies have masked key information within specific datasets, including input questions in NLP benchmarks Deng et al. (2024b); Golchin and Surdeanu (2023b); Liu et al. (2024), column names in SQL code generation questions Ranaldi et al. (2024), character names in books Chang et al. (2023), and labels in NLI and SST tasks Magar and Schwartz (2022).
@@ 4.3.3 Generation Variation
This type of approach explores how an LM’s outputs vary if it has seen an instance during training.
@@
$\text{Var}(\{M_{\cdot}(x_{p})\})\left\{\begin{array}[]{ll}<\xi_{v}&\text{if }f(M,x_{p})=1\\ \geq\xi_{v}&\text{if }f(M,x_{p})=0\end{array}\right.$ (7)
where $Var(\{M_{\cdot}(x_{p})\})$ measures the variations of outputs from $M$ under diverse, different sampling strategies when given $x_{p}$; $\xi_{v}$ is a threshold, based on the type of input $x_{p}$ and sampling strategies.
@@ 4.3.4 Metadata-based Memorization
$\left\{\begin{array}[]{ll}\exists x\in D,S(M(x_{m}),x)\geq\tau_{m}&\text{if $D$ is seen}\\ \forall x\in D,S(M(x_{m}),x)<\tau_{m}&\text{if $D$ is unseen}\end{array}\right.$ (8)
Here, $M(x_{m})$ is the set of instances that $M$ generates when given $D$’s metadata $m$; $S(M(x_{m}),x)$ represents the highest similarity between $x$ and any instance $x^{\prime}$ as a subsequence of $M(x_{m})$; $\tau_{m}$ is the contamination threshold for $S(\cdot,\cdot)$.
@@ 4.4 Answer Memorization
Answer memorization is usually conducted at the dataset level. It introduces perturbations to the original dataset, measures the LM’s performance change, and aims to detect whether the LM’s high performance is due to memorizing its answer.
@@
Given an LM $M$ and an evaluation dataset $D$, one can generate a similar dataset $D^{\prime}$ that is unseen by $M$, by modifying every $x\in D$.
@@
$\text{Eval}(M,D)\left\{\begin{array}[]{ll}\gg\text{Eval}(M,D^{\prime})&\text{if $D$ is seen}\\ \not\gg\text{Eval}(M,D^{\prime})&\text{if $D$ is unseen}\end{array}\right.$ (9)